# SCHEME: Scalable Channer Mixer for Vision Transformers

### v1

@@ 3.2 The SCHEMEformer family
The proposed SCHEME module can be plugged into any ViT architecture, enabling efficient control of model complexity via the mixer hyperparameters $g_{1}$, $g_{2}$, and $E$. Table 1 shows that naively scaling down the ViT model by simply reducing $E$ causes a significant accuracy loss. The SCHEME module allows much more effective control of the accuracy/complexity trade-off, producing models of better performance for a fixed computational budget. This is demonstrated by the introduction of a new family of models, denoted as SCHEMEformer, obtained by replacing the channel mixer of the Metaformer-PPAA [57] architecture with the SCHEME module. Two such configurations are shown in Table 2 where the naming follows the convention {model-name}-{$g_{1}$$g_{2}$}-e{$E$}. These are shown to achieve SOTA accuracy for lower computational budgets in the experiments section, where we also discuss results with other backbones.
S12 S24 S36
P F P F P F
Model (M) (G) (M) (G) (M) (G)
SCHEMEformer-PPAA-44-e8 12 1.8 21 3.3 55 7.96
SCHEMEformer-PPAA-12-e8 21 3.3 40 6.5 59 9.6
Table 2: Scaling MetaFormer-PPAA using SCHEME. Each column of the table reports to one baseline model (S12 to S36) of increasing size. P and F denote Parameters and FLOPs respectively. SCHEMEformer-PPAA-44-e8 (SCHEMEformer-PPAA-12-e8) models use configurations $g_{1}$, $g_{2}$, and $E$ that downscale (upscale) the baseline.
3.3 Computational Complexity
The complexity of the BD-MLP block is controlled by the group numbers $g_{1}$, $g_{2}$ and the expansion factor $E$, with a total cost $\mathcal{O}(Ed^{2}/g_{1}+Ed^{2}/g_{2})$, where $d$ is the channel dimension. The computational cost of CCA is $\mathcal{O}(Nd^{2})$ where $N$ is the number of tokens. Since CCA is not used during inference, it only adds to the computations during training. The SCHEME framework provides a systematic way to control the trade-off of transformer width vs depth, by controlling the block size and expansion hyperparameters.
@@ 4 Experimental Results
In this section, we discuss an evaluation of the SCHEME module on Image Classification (ImageNet-1K [11]), Object Detection (COCO-17 [28]), and Semantic Segmentation (ADE-20K [65]) benchmarks.
#Params FLOPs Top-1 Acc
Model  (M) (G) (%)
Attention Mixers
Poolformer-S7 [57] 8.60 1.1 73.0
PVT-Tiny [51] 13 1.9 75.1
XCiT-T12 [16] 7 1.2 77.1
Poolformer-S12 [57] 12 1.8 77.2
MetaFormer-PPAA-11-e2-S12 [57] 12 1.8 78.9
XCiT-T24 [16] 12 2.3 79.4
ViT-B/16 [15] 86 17.6 79.7
SCHEME-Poolformer-44-e8-S12 7.20 1.0 73.0
SCHEME-Poolformer-12-e8-S12 16.7 2.6 78.5
SCHEME-CSWin-44-e8-Tiny 15.6 3.0 79.2
SCHEME-Swin-44-e8-Tiny 19.7 3.1 79.6
SCHEMEformer-PPAA-44-e8-S12 12 1.77 79.7
PVT-Small [51] 25 3.8 79.8
QuadTree-B-b1 [41] 14 2.3 80.0
Poolformer-S24 [57] 21.0 3.4 80.3
MetaFormer-PPAA-S12 [57] 17 2.6 81.0
PVT-Medium [51] 44 6.7 81.2
Swin-Tiny [32] 29 4.5 81.3
T2T-ViT t-14 [60] 22 6.1 81.7
DeiT-B [43] 86 17.5 81.8
CaiT-XS-24 [45] 27 5.4 81.8
ViL-Small [61] 25 5.1 82.0
XCiT-S12 [16] 26 4.8 82.0
SCHEME-Poolformer-12-e8-S24 30.8 4.9 80.5
SCHEME-T2T-ViT-44-e8-14 17.4 6.1 81.7
SCHEME-Swin-12-e8-Tiny 36.9 5.9 81.7
SCHEMEformer-PPAA-44-e8-S24 21 3.31 81.9
SCHEMEformer-PPAA-12-e8-S12 21 3.35 82.0
SCHEME-T2T-ViT-12-e8-14 27.7 8.1 82.1
SCHEME-DaViT-44-e8-Tiny 19.7 3.8 82.1
Focal-Tiny [56] 29 4.9 82.2
CrossViT-Base [2] 105 21.2 82.2
T2T-ViT-24[60] 64 14.1 82.3
CPVT-Base [7] 88 17.6 82.3
DaViT-Tiny [12] 23.0 4.3 82.8
CSWin-Tiny [13] 23.0 4.3 82.8
SCHEMEformer-PPAA-12-e8-S24 40 6.47 82.8
SCHEME-DaViT-12-e8-Tiny 37.0 6.6 83.0
SCHEME-CSWin-12-e8-Tiny 29.1 5.6 83.2
XCiT-L24 [16] 189 36.1 82.9
TNT-B [20] 66 14.1 82.9
Swin-Small [32] 50 8.7 83.0
ViL-Base [61] 56 13.4 83.2
CSWin-Small [13] 35.0 6.9 83.6
PVTv2-B5 [52] 82.0 11.8 83.8
Focal-Base [56] 89.8 16.0 83.8
SCHEMEformer-PPAA-44-e8-S36 55 7.96 83.2
SCHEMEformer-PPAA-12-e8-S36 58 9.58 84.0
Convolution+Attention Mixers
CeiT-S [59] 24 4.8 82.0
CvT-13 [53] 20 4.5 81.6
CoAtNet-0 [9] 25 4.2 81.6
PatchConv-S60 [44] 25 4.0 82.1
CvT-21 [53] 32 7.1 82.5
ViTAEv2-S [63] 19 5.7 82.6
$*{}^{*}$ 17.8 3.4 80.7
$*{}^{*}$ 24 4.1 81.6
SCHEME-CAformer-44-e8-S18 19 2.92 82.4
SCHEME-CAformer-12-e8-S12 24 3.61 82.9
Table 3: Image Classification on ImageNet-1K dataset. Comparison with SOTA ViT models grouped by the type of token mixers and accuracy. Proposed SCHEMEformer-PPAA-44-e8(12-e8) models are shown in light blue and the baseline architectures are shown in light cyan. *{*}* denotes our implementation as their code is not publicly available.
Figure 3: Comparison of various SCHEMEformer (blue stars) and SOTA ViT architectures (orange circles) (calculated on NVIDIA-Titan-X GPU with batch size 1). Left: Accuracy-Throughput trade-off. Right: Accuracy-Parameters trade-off.
@@ 4.1 Comparisons to the state of the art
Implementation Details: For image classification, we use a batch size of 1024, adjusted according to memory availability, and train all models on 8 GPUs with AdamW optimizer and cosine annealing learning rate schedule, of learning rate 1e-3. Warmup learning is applied in the first 5 epochs. Object detection is based on the setting of [57], where all backbones are trained with RetinaNet [29] and Mask-RCNN [22] detection heads, for a fair comparison with prior methods. A 1x learning schedule (12 epochs with AdamW optimizer and initial learning rate of 1e-4) is used and all models are implemented with the mmdetection [3] codebase, capping input sizes at 800, 1333 for the shorter and longer sides, respectively. For semantic segmentation, following prior methods [4, 24, 57], we train with a batch size of 16 for 80,000 iterations with AdamW optimizer and an initial learning rate of 2e-4. We use the mmsegmentation [8] codebase with an input size $512\times 512$ for training and resize the shorter side to 512 pixels for testing. For both detection and segmentation, imagenet pretrained models are used to initialize the backbones.
#Par $AP$ $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$ RetinaNet 1× ResNet-18[21] 21.3 31.8 49.6 33.6 16.3 34.3 43.2 PoolFormer-S12[57] 21.7 36.2 56.2 38.2 20.8 39.1 48.0 ResNet-50[21] 37.7 36.3 55.3 38.6 19.3 40.0 48.8 SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 PoolFormer-S24[57] 31.1 38.9 59.7 41.3 23.3 42.1 51.8 ResNet-101[21] 56.7 38.5 57.8 41.2 21.4 42.6 51.1 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 Mask R-CNN 1× #Par $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ Mask R-CNN 1× ResNet-18[21] 31.2 34.0 54.0 36.7 31.2 51.0 32.7 PoolFormer-S12[57] 31.6 37.3 59.0 40.1 34.6 55.8 36.9 ResNet-50[21] 44.2 38.0 58.6 41.4 34.4 55.1 36.7 SCHEMEformer-44-e8-S12 31 39.8 61.9 42.9 24.1 53.0 42.3 PoolFormer-S24[57] 41.0 40.1 62.2 43.4 37.0 59.1 39.6 ResNet-101[21] 63.2 40.4 61.1 44.2 36.4 57.7 38.8 SCHEMEformer-44-e8-S24 41 40.9 62.5 44.6 24.6 55.6 43.8 Table 4: COCO-17 Object Detection and Instance Segmentation. All backbones are pretrained on ImageNet-1K and use 1x learning schedule. $AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively.
Backbone #Par FLOPs mIoU (%) ResNet-18[21] 16 32.2 32.9 PVT-Tiny[51] 17 33.2 35.7 ResNet-50[21] 29 45.6 36.7 PoolFormer-S12[57] 16 30.9 37.2 ResNet-101[21] 48 65.1 38.8 ResNeXt-101-32x4d[55] 47 64.7 39.7 PVT-Small[51] 28 44.5 39.8 XCiT-T12/8[16] 8.4 - 39.9 ResNeXt-101-64x4d[55] 86 103.9 40.2 PoolFormer-S24[57] 23 39.3 40.3 SCHEMEformer-44-e8-S12 15.5 34.3 40.9 PVT-Medium[51] 48 61.0 41.6 PoolFormer-S36[57] 35 47.5 42.0 PVT-Large[51] 65 79.6 42.1 PoolFormer-M36[57] 60 67.6 42.4 SCHEMEformer-44-e8-S24 24.8 45.7 42.5 Table 5: Semantic Segmentation results on ADE20K dataset using Semantic FPN [24]. All backbones are pretrained on ImageNet-1K. FLOPs are calculated at $512\times 512$ resolution.
Model BDM CCA Acc (%) Baseline 78.9 44-e8-S12 79.1 (+0.2) 44-e8-S12 79.7 (+0.6) Table 6: Ablation study of the contribution of BD-MLP and CCA branches to the overall performance for SCHEMEformer-PPAA-44-e8-S12 model. Baseline refers to Metaformer-PPAA-11-e2-S12. Module #Par FLOPs Top-1  (M) (G) Acc (%) Shuffle 11.83 1.77 79.1 SE 11.98 1.77 79.3 Conv 13.31 1.97 79.6 DyCCA 11.90 1.83 79.6 CCA 11.83 1.77 79.7 Table 7: Ablation study of alternative designs for feature attention branch in SCHEME.
@@
Image Classification: Image classification is evaluated on Imagenet-1K, without using extra data. We report the results with single crop top-1 accuracy at $224\times 224$ input resolution. Results are presented for various backbones and hyperparameter configurations, to show the generalizability of the SCHEME module. We evaluate the SCHEMEformer family of models based on the Metaformer-PPAA-S12 [57] obtained by replacing the MLP of the original model with the SCHEME module. Additionally, we replace the pooling of the first two stages of the Metaformer-PPAA-S12 with convolution  [58] to evaluate SCHEME-CAformer models.
Table 3 compares SCHEME models to various SOTA models from the literature grouped by the type of token mixers and accuracy. The SCHEME models are highlighted in blue and the baseline models on which they are based in cyan. While the table is ordered by increasing accuracy, meaningful comparisons must account for both accuracy and parameter size or FLOPs, since the models have very different sizes. The SCHEMEformer family achieves the best trade-off between accuracy and computation or parameter size by using a larger internal feature dimension for the MLP where the gains are larger for smaller models, whose performance is not near saturation. For example, as highlighted in red font, for the same compute and parameter cost, the SCHEMEformer-PPAA-44-e8-S12 improves on the performance of the baseline MetaFormer-PPAA-11-e2-S12 by +0.8% and achieves a SOTA result of 79.7% top-1 accuracy at 1.77G FLOPs. The table also shows the accuracy of the SCHEME module applied to other popular ViT architectures, such as T2T-ViT, Swin, CSWin, Da-ViT in the Attention Mixers category, and CoatNet, CAformer in the Convolution+Attention Mixers category realizing models of intermediate complexity with comparable or better accuracy for the same parameters and FLOPs. For example, SCHEME-CoAtNet-12-e8-0 achieves 81.6% accuracy with 1M fewer parameters and 0.1G fewer FLOPs than the baseline CoAtNet-0 (highlighted in brown). SCHEME-CSWin-12-e8-Tiny (blue) achieves 83.2% accuracy with just 29 M parameters and 5.6 G FLOPs sandwiched between CSWin-Tiny (82.8%) and CSWin-Small (83.6%).
Figure 1 provides a visual illustration of the results in the table. The plot on the left compares the SCHEME family (blue stars) to all other models (orange dots) in terms of the trade-off between FLOPS and accuracy. It is clear that most SCHEME models lie on the Pareto frontier for these two objectives, significantly expanding the set of models on this frontier. This illustrates the fine control that SCHEME allows over the accuracy/complexity trade-off of transformer models. The plot on the right shows how the SCHEMEFormer models improve on the accuracy/complexity trade-off of the MetaFormer model on which they are based. This shows that the SCHEME module generalizes well to different backbones and offers higher flexibility in scaling.
Throughput and Memory: Beyond higher accuracy, SCHEMEformer models also have low latency and parameter counts. The left of Figure 3 shows a favorable throughput vs accuracy trade-off for various SCHEME models compared to SOTA ViT architectures. Throughput is computed on NVIDIA-Titan-X GPU with a batch size of 1 averaged over 1000 runs. Despite having larger internal dimensions, the SCHEMEformer-PPAA-12-e8-S12 has a throughput (130 images per second) comparable to that of the baseline MetaFormer-PPAA-11-e2-S12. This is a consequence of the block diagonal structure, which enables higher FLOPs utilization [18]. The right of Figure 3 shows a favourable accuracy vs parameters tradeoff comparison for SCHEME models compared to SOTA ViT architectures. SCHEME based scaling achieves higher accuracy with lower parameter counts. These results show that SCHEME models lie on the Pareto frontiers of accuracy vs throughput and vs memory consumption.
Object Detection: Table 4 compares SCHEMEformer models to models of similar complexity on the COCO-17 object detection and instance segmentation benchmark, for both Retinanet and Mask-RCNN detection heads. The table is ordered by increasing mAP ($AP$, $AP^{b}$) values. Again, the SCHEMEformer models outperform all other models of the same or smaller size. For example, SCHEMEformer-44-e8-S12 with RetinaNet head achieves 38.3% overall mAP with just 21M params, outperforming models of comparable size by more than 2% and achieving comparable performance to models with $2.5$ times their size. Similar gains are obtained for the Mask-RCNN based SCHEMEformer-44-e8-S12. These results show that the SCHEME module also advances object detection.
Semantic Segmentation: Table 5 compares SCHEMEformer-44-e8 S12 and S24 models using the semantic FPN framework [24] to other models of similar complexity on ADE20K semantic segmentation benchmark. As before, the SCHEMEformer models outperform all other models of the same or smaller size. For example, SCHEMEformer-44-e8-S12 achieves 40.9% mIoU, which is 5.2% higher than the model of most comparable size, PvT-Tiny. Similarly, the S24 model outperforms PoolFormer-M36 using only 41% of its parameters. In general, the gains obtained for semantic segmentation are higher than those of object detection, suggesting that the feature clusters formed by the block diagonal structure have larger benefits for dense prediction tasks.

### v4

@@ 3.2 The SCHEMEformer family
The proposed SCHEME module enables efficient control of model complexity via the mixer hyperparameters $g_{1}$, $g_{2}$, and $E$. Table 3 shows that naively scaling down the ViT model by simply reducing $E$ causes a significant accuracy loss. The SCHEME module allows much more effective control of the accuracy/complexity trade-off, producing models of better performance for a fixed computational budget. This is demonstrated by the introduction of a new family of models, denoted as SCHEMEformer, obtained by replacing the channel mixer of the Metaformer-PPAA [55] architecture with the SCHEME module. See Table 3 for two such configurations.
@@ 4 Experimental Results
Figure 3: Impact of CCA (SCHEMEformer-44-e8-S12). Left: Evolution of weight $1-\alpha$ across model layers. Right: Class separability of output features $\bf y$ (over 50 random classes of ImageNet-1K validation set) for model trained with and without CCA. See Appendix Figure 11 for zoomed version.
Figure 4: SCHEME tradeoffs. Left: Accuracy vs FLOPs of various SCHEME models with different MLP configurations. Right: SCHEME mixer improves accuracy for fixed throughput or vice-versa for various popular ViT architectures. See Appendix Figure 11 for zoomed version.
@@ 4.1 Comparisons to the state of the art
Image Classification: Image classification is evaluated on Imagenet-1K, without using extra data. We report the results with single crop top-1 accuracy at $224\times 224$ input resolution. We evaluate the SCHEMEformer family of models based on the Metaformer-PPAA-S12 [55] obtained by replacing the MLP of the latter with the SCHEME module. Refer to Appendix A.2 for implementation details.
We start by evaluating how this improves the trade-off between model accuracy and complexity. As discussed in Section 3.1, when $g_{1}=g_{2}$, a SCHEME transformer of expansion factor $gE$ has identical complexity to a standard transformer of expansion factor $E$. Hence, for fixed FLOPS, SCHEME allows an increase of the expansion factor by $g$. Figure 4 a) compares the performance of the Metaformer-PPAA-S12 with expansion ratios $E\in\{1,2,4,8\}$ to comparable variants of the SCHEMEformer-PPAA-S12, with SCHEME mixers of either $g=2$ (green curve) or $g=4$ (blue curve) groups. The SCHEMEformer models have a better trade-off between accuracy and FLOPS, achieving higher accuracies for all complexity levels. Among these, the one with more feature groups ($g=4$) has the best performance. While SCHEMEformer gains are observed for all FLOP levels, they are larger for lower complexity models. This makes SCHEME particularly attractive for the design of low complexity transformers, e.g. for edge devices.
We next compare the SCHEMEformer family against the SOTA transformers in the literature. This is not an easy comparison, since models vary in size, FLOPS, and throughput. Because it is difficult to make any of these variables exactly the same for two different architectures, the comparison is only possible in terms of how the different architectures trade-off accuracy for any of the other factors. For a given pair of variables, e.g. FLOPS vs accuracy, the model is said to be on the Pareto frontier of the two variables if it achieves the best trade-off between the two. Table 4.1 presents a comparison of the SCHEMEformer family against various SOTA transformers in the literature. Each section of the table compares a SCHEMEformer model to a group of SOTA transformers of equivalent size or complexity. Note that, in each section, the remaining models have both lower throughput and accuracy than the SCHEMEformer model. In many cases they also have more parameters and FLOPs. Figure 1 provides a broader visualization of how SCHEME models establish Pareto frontiers for accuracy vs FLOPS, accuracy vs throughput, and accuracy vs model size (parameters). Figure 1 a) illustrates the trade-off between accuracy and FLOPS of many SOTA transformers. The dashed line connects the SCHEMEformer model results, summarizing the accuracy-FLOPs trade-off of the family. It can be seen that the SCHEME models lie on the Pareto frontier for these two objectives. This illustrates the fine control that SCHEME allows over the accuracy/complexity trade-off of transformer models. Figure 1 b) presents a similar comparison for model sizes. Like for FLOPS, the SCHEME models lie on the Pareto frontier for accuracy vs model size. In fact the two plots are quite similar, showing that in general there is a good correlation between model size and FLOPS.
Table 4: Image Classification on ImageNet-1K. Comparison of SCHEME models using expansion ratio 8 with SOTA ViTs grouped by accuracy. SCHEME family has higher throughput and accuracy than SOTA models.
Model #Params FLOPs Thru Top-1 Acc
 (M) (G) (im/s) (%)
gMLP-Ti [25] 6 1.4 - 72.3
ViT-L/16 [11] 307 63.6 37 76.1
Meta-11-e2-S12 [55] 12 1.8 133 78.9
MogaNet-T [23] 5 1.10 44 79.0
XCiT-T24 [12] 12 2.3 - 79.4
ViT-B/16 [11] 86 17.6 112 79.7
SCHEME-44-S12 12 1.77 133 79.7
$S^{2}$-MLP-deep [54] 51 10.5 - 80.7
Meta-S12 [55] 17 2.6 87 81.0
Swin-Tiny [27] 29 4.5 100 81.3
T2T-ViT t-14 [57] 22 6.1 70 81.7
DeiT-B [37] 86 17.5 114 81.8
ViL-Small [59] 25 5.1 - 82.0
SCHEME-12-S12 21 3.35 130 82.0
Focal-Tiny [53] 29 4.9 29 82.2
CPVT-Base [4] 88 17.6 - 82.3
DaViT-Tiny [8] 23.0 4.3 61 82.8
CSWin-Tiny [9] 23.0 4.3 20 82.8
SCHEME-12-S24 40 6.47 69 82.8
XCiT-L24 [12] 189 36.1 - 82.9
Swin-Small [27] 50 8.7 31 83.0
ViL-Base [59] 56 13.4 - 83.2
CSWin-Small [9] 35 6.9 11 83.6
SCHEME-12-S36 58 9.58 38 84.0
Table 5: Semantic Segmentation results on ADE20K. FLOPs calculated at $512$ resolution for Semantic FPN and $1024$ resolution for UperNet.
Backbone #P F mIoU
(M) (G) (%)
Semantic FPN
ResNet-18[19] 16 32.2 32.9
PVT-Tiny[44] 17 33.2 35.7
ResNet-50[19] 29 45.6 36.7
PoolFormer-S12[55] 16 30.9 37.2
ResNet-101[19] 48 65.1 38.8
ResNeXt-101-32x4d[51] 47 64.7 39.7
PVT-Small[44] 28 44.5 39.8
XCiT-T12/8[12] 8.4 - 39.9
PoolFormer-S24[55] 23 39.3 40.3
SCHEMEformer-44-S12 15.5 34.3 40.9
PVT-Medium[44] 48 61.0 41.6
PoolFormer-S36[55] 35 47.5 42.0
PVT-Large[44] 65 79.6 42.1
PoolFormer-M36[55] 60 67.6 42.4
SCHEMEformer-44-S24 24.8 45.7 42.5
UperNet
Swin-Tiny [27] 60 945 44.5
PVT-Large [45] 65 318 44.8
Focal-Tiny [53] 62 998 45.8
XCiT-S12/16 [12] 52 45.9
DaViT-Tiny [8] 60 940 46.3
SCHEME-DaViT-12-Tiny 68 969 47.1
This is not the case for throughput, which for transformers is known to not necessarily correlate with FLOPS, due to GPU parallelism. For example, Table 4.1 shows that, CSWin models have lower throughput despite having lower FLOPs while ViT has higher throughput despite having higher FLOPs. SCHEMEformer controls this trade-off by controlling the expansion ratio and block diagonal structure, which enables higher FLOPs utilization for a given throughput [16]. We compute throughput on a NVIDIA-Titan-X GPU with a batch size of 1 with input size 224x224 averaged over 1000 runs. While comparisons could be made for larger batch sizes, we consider the setting for live/streaming applications, where speed is most critical. In these applications, the concern is usually inference throughput, which requires batch size of 1. Figure 1 c) illustrates the trade-off between accuracy and throughput of various models, including very fast ResNet models of low accuracy. It can bee seen that the SCHEMEformer again achieves the best trade-off between these two variables, thus lying on the Pareto frontier for accuracy vs throughput. Its performance is particularly dominant in the range of throughputs between 75 and 150 images/sec, where it significantly outperforms the other methods. These results demonstrate how the SCHEME module endows transformer designers with the ability to produce multiple models at different points of the Pareto frontiers of accuracy vs model size, FLOPS, or throughput.
SCHEME with other ViTs: Figure 4 shows the Accuracy-Throughput curves for various SCHEME models obtained by replacing the MLP blocks of popular ViT architectures. Table A in Appendix shows the detailed accuracy-latency comparisons of SCHEME with recent ViTs. The table shows the results for 12 different ViT backbones ranging from small (7.5 M) to large (143 M) models. It shows that SCHEME outperforms the baseline ViT models across on all 12 models either in accuracy or throughput. This demonstrates that SCHEME improves the accuracy for a fixed throughput or the throughput for a fixed accuracy, or both. The gains can be substantial, e.g. about 1.5% accuracy gain (constant throughput) for the fastest transformer or speed gains of up to 20% at constant accuracy. The table also shows that SCHEME is effective when scaling the model size as seen with SCHEME-ViT Large model. This shows that SCHEME benefits various ViT backbones and not just the MetaFormer.
Semantic Segmentation: Table 4.1 compares the semantic segmentation performance of two SCHEMEformer models using the semantic FPN framework [21] to various SOTA models of similar complexity, on ADE20K. Since we do not not have access to the throughput of most models, we report only parameter sizes and FLOPS. In each section of the table, the remaining models have comparable or larger FLOPs and model size but lower accuracy than the corresponding SCHEMEformer. For example, SCHEMEformer-44-e8-S12 achieves 40.9% mIoU, which is 5.2/3.7 points higher than the PvT-Tiny/PoolFormer-S12, which both have comparable size and FLOPs. Similarly, the S24 model outperforms PoolFormer-M36 using only 41% of its parameters. To demonstrate the applicability of SCHEME to larger models, we also present a comparison of the SCHEME version of the DaViT-Tiny using the UperNet framework [50]. While the DaViT-Tiny is already the best model in the table, the use of the SCHEME mixer improves its performance by an additional 0.8 points.
Object Detection: Table 5 compares SCHEMEformer models to models of similar complexity on the COCO-17 object detection and instance segmentation benchmark, for both Retinanet and Mask-RCNN detection heads. Again, the SCHEMEformer models outperform most other models of the same or smaller size. The only exception is the PoolFormer-S24, which slightly outperforms ($0.1$ points) the comparable SCHEME-former-44-S24, for the RetinaNet head. However, with the stronger Mask R-CNN head, the SCHEMEformer-44-S24 beats the PoolFormer-S24 by $0.8$ points. For the top performing models in the bottom third of each section of Table 5, SCHEME-DaViT-Tiny improves over DaViT-Tiny by an additional 0.7% and 0.9%, for RetinaNet and MaskRCNN heads respectively, while maintaining a comparable throughput.
Table 6: COCO-17 Object Detection and Instance Segmentation. All backbones pretrained on ImageNet-1K (1x learning schedule). ($AP^{b}$, $AP^{m}$): (bounding box AP, mask AP). $T$: throughput (images/sec).
RetinaNet 1× #Par T $AP$ $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$
PoolFormer-S12[55] 21.7 13.1 36.2 56.2 38.2 20.8 39.1 48.0
ResNet-50[19] 37.7 15.7 36.3 55.3 38.6 19.3 40.0 48.8
SCHEME-44-e8-S12 21 6.1 38.3 58.0 40.4 21.0 41.4 52.3
PoolFormer-S24[55] 31.1 8.9 38.9 59.7 41.3 23.3 42.1 51.8
ResNet-101[19] 56.7 12.1 38.5 57.8 41.2 21.4 42.6 51.1
SCHEME-44-e8-S24 31 3.5 38.8 58.7 41.2 22.5 41.5 53.5
DAT-T [49] 38 - 42.8 64.4 45.2 28.0 45.8 57.8
CrossFormer-S [46] 41 - 44.4 55.3 38.6 19.3 40.0 48.8
DaViT-Tiny [8] 39 8.2 44.0 65.6 47.3 29.6 47.9 57.3
SCHEME-DaViT 47 7.8 44.7 66.2 48.3 30.0 48.8 57.2
Mask R-CNN 1× #Par T $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
PoolFormer-S12[55] 31.6 9.9 37.3 59.0 40.1 34.6 55.8 36.9
ResNet-50[19] 44.2 15.4 38.0 58.6 41.4 34.4 55.1 36.7
SCHEME-44-e8-S12 31 6.0 39.8 61.9 42.9 37.1 59.2 39.4
PoolFormer-S24[55] 41.0 7.9 40.1 62.2 43.4 37.0 59.1 39.6
ResNet-101[19] 63.2 12.1 40.4 61.1 44.2 36.4 57.7 38.8
SCHEME-44-e8-S24 41 3.4 40.9 62.5 44.6 37.8 59.7 40.4
DAT-T [49] 48 - 44.4 67.6 48.5 40.4 64.2 43.1
CrossFormer-S [46] 50 - 45.4 68.0 49.7 41.4 64.8 44.6
DaViT-Tiny [8] 48 7.8 45.0 68.1 49.4 41.1 64.9 44.2
SCHEME-DaViT 57 7.4 45.9 68.3 50.2 41.5 65.4 44.3
Table 7: Contribution of BD-MLP and CCA branches in SCHEME.
Model BDM CCA Acc (%)
Baseline 78.9
44-e8-S12 79.1
44-e8-S12 79.7
Table 8: Alternative feature attention designs in SCHEME.
Module #Par FLOPs T Top-1
 (M) (G) (img/s) Acc (%)
Shuffle 11.83 1.77 108 79.1
SE 11.98 1.77 86 79.3
Conv 13.31 1.97 95 79.6
DyCCA 11.90 1.83 75 79.6
CCA 11.83 1.77 133 79.7
Figure 5: Ablation study on the formation of feature clusters in the BD-MLP branch of the SCHEME module.
CCA G1 G2 G3 G4 Ensemble
51.1 54.3 55.4 54.7 73.2
47.6 50.0 49.2 55.1 73.8